MC ATC AVC P=150 D=MR 120 104 1500 2400 3200 4800 6000 7400
a-what is the firm's maximizing output?
b-what is the firm's profit?
c- if the price drops below ? this firm will shut down
Answer #1

a. Firm’s maximizes output at the point where P=MR=MC. At this point, equilibrium quantity is 6000 units and equilibrium price is $150.

b. Profit = total revenue - total cost = P*Q - ATC*Q = 150*6000 - 120*6000 = $ 180,000

c. If the price drops below AVC (i.e, P= $ 66) then the firm will shut down.